Solve each of the following systems by addition. If a unique solution does not exist, state
whether the system is inconsistent or dependent.
2x + 3y = 1
5x + 3y = 16

Answer by jainenderkapoor(61)


2x + 3y = 1 ---- (1)
5x + 3y = 16 ---- (2)
Subracting (1) from (2) we get
5x + 3y - 2x - 3y = 16 - 1
5x - 2x + 3y - 3y = 16 - 1
3x = 15
3x/3 = 15/3 (Dividing both sides by 3)
x = 5
Putting this value of x in (1) we get
2(5) + 3y = 1
10 + 3y = 1
10 +3y - 10 = 1-10 (Subtracting 10 from both sides)
3y = -9
3y/3 = -9/3 (Dividing both sides by 3)
y = -3
So the unique solution is (5, -3)
Hope the solution is clear to you. In case you have any doubt, you are welcome to contact me at jainenderkapoor@gmail.com. I am also available for online tutoring at reasonable rates.